# Dynamics of Meta-learning Representation in the Teacher-student Scenario

### v1

@@ Abstract
Gradient-based meta-learning algorithms have gained popularity for their ability to train models on new tasks using limited data. Empirical observations indicate that such algorithms are able to learn a shared representation across tasks, which is regarded as a key factor in their success. However, the in-depth theoretical understanding of the learning dynamics and the origin of the shared representation remains underdeveloped. In this work, we investigate the meta-learning dynamics of the non-linear two-layer neural networks trained on streaming tasks in the teach-student scenario. Through the lens of statistical physics analysis, we characterize the macroscopic behavior of the meta-training processes, the formation of the shared representation, and the generalization ability of the model on new tasks. The analysis also points to the importance of the choice of certain hyper-parameters of the learning algorithms.
@@ I Introduction
Supervised learning has long been the dominant paradigm in machine learning and artificial intelligence, yielding substantial achievements across various fields [1]. The primary objective of supervised learning is to obtain a function mapping tailored to a specific task by training the machines on a set of input-output examples. However, one major limitation of this paradigm is that the trained machine is adapted to the specific task considered, and is typically not suited for other tasks, even though they may share some common properties with the existing task. In extreme scenarios, it may require training a new model from scratch for each new task. This is not desirable if the supervised learning tasks require a large number of input-output examples, a common requirement for training large-scale deep neural networks [2, 3].
@@
To overcome this drawback, many alternative machine learning pipelines have been proposed, such as multitask learning, transfer learning, self-supervised learning, meta-learning, and so on [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A common strategy in the recent deep learning practice is to learn a useful representation of the data, which will become applicable to many different downstream tasks [6, 7, 8]. Alternatively, meta-learning typically aims at learning to how learn the algorithms for new tasks. Among various meta-learning approaches, the gradient-based meta-learning algorithms have received a lot of attention [11, 12]. A notable example is the Model-Agnostic Meta-Learning (MAML) algorithm [12]. It aims to learn a good initialization of the model parameters such that the model can be quickly adapted to new tasks starting from this set of initial model parameters; these initial model parameters are then meta-trained on many different tasks. As such, it involves two nested levels of optimization processes, both of which are implemented by gradient descent algorithms in MAML. Subsequent empirical work speculated that the success of MAML is also achieved by learning a shared representation across tasks and further proposed methods to simplify MAML [14].
@@
However, the nested optimization processes make it challenging to develop theoretical analyses for meta-learning algorithms like MAML. To facilitate the analysis, researchers often utilize simplified architectures, such as linear models [15, 16]. Notably, recent works investigated the learning dynamics of two-layer linear networks and proved that they can successfully learn the hidden representation under the learning dynamics with the population gradients in use [17, 18]. It is unknown whether similar results can be extended to more realistic settings of non-linear neural networks and the learning dynamics with finite-sample gradients. In this paper, we aim to tackle the challenging scenarios with non-linear network by employing tools from statistical physics, which has a longstanding tradition in the theoretical development of artificial neural networks [19, 20, 21], and has regained momentum in recent years [22, 23]. These methods have been recently applied to the multitask learning scenario [24, 25, 26, 27, 28] to study phenomena such as catastrophic forgetting and the effects of task similarity, as well as to propose novel learning algorithms; in these studies, the number of tasks is finite and there are abundant examples for each task. In this work, we focus on a scenario that is more typical in meta-learning, where the number of tasks is large, but the number of examples per task is much smaller than in single-task settings. We will further consider that the tasks arrive in a streaming, online fashion, which is similar to the usual continual learning setting. However, meta-learning approaches have greater emphasis on the ability to adapt to new tasks and require additional training steps for a new task, which differs from the usual continual learning setting.
@@ II The Problem and Algorithm of Meta-learning
The primary objective of meta-learning is to create a framework where a model can learn to adapt to some new task using a few examples. The model involves a meta-learner which is trained on different tasks in the meta-training phase, aiming to develop a meta-level knowledge that are useful for future tasks.
@@
In the meta-learning problem setup, there are typically many tasks available, which are denoted as $\{\mathcal{T}_{t}\}_{t=1,2,\dots}$ and assumed to be drawn from a certain distribution $\mathcal{P}(\mathcal{T})$. Each task $\mathcal{T}_{t}$ comes with its own training and validation datasets, denoted as $\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}}$ and $\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}$, respectively. We focus on supervised learning tasks, so both the training and validation sets consist of input-output example pairs. Suppose a specific learner $\mathcal{L}_{t}$ is assigned to solve task $\mathcal{T}_{t}$. Its model weights are trained on the training set $\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}}$ and evaluated on the validation set $\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}$. If the learner $\mathcal{L}_{t}$ inherits useful meta-level knowledge about the task $\mathcal{T}_{t}$ from an experienced meta-learner, it should require significantly fewer training examples than if it were learning from scratch. The challenge is how to effectively obtain such an experienced meta-learner. The Model-Agnostic Meta-Learning (MAML) algorithm addresses this by maintaining and adapting the meta-learner which serves as a valuable initialization (denoted as $\boldsymbol{\theta}_{0}$) for the model weights $\boldsymbol{\theta}_{t}$ of the learner $\mathcal{L}_{t}$. Since the learner $\mathcal{L}_{t}$ does not start from scratch, but has inherited the good initialization $\boldsymbol{\theta}_{0}$ from the meta-learner, it may require only a few steps of gradient-descent update by using the training set $\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}}$ in order to solve the task $\mathcal{T}_{t}$. The parameters $\boldsymbol{\theta}_{0}$ of the meta-learner are then updated also in a gradient-descent manner in order to reduce the validation error of the tasks that have shown up. Note that the validation error of task $\mathcal{T}_{t}$ is defined on the validation set $\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}$.
@@
Formally, the learner $\mathcal{L}_{t}$ would like to learn a function $f_{\boldsymbol{\theta}_{t}}$ parameterized by $\boldsymbol{\theta}_{t}$ for solving task $\mathcal{T}_{t}$. It does so by starting from the meta-learner’s weights $\boldsymbol{\theta}_{0}$ and performing gradient descent using the training data $\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\mu},\sigma^{t,\mu})\}_{\mu=1}^{P}$. In MAML, a one-step gradient descent is applied [12],
$\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{0}-\eta_{t}\big{[}\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}\mid\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (1)
@@
$\ell(\boldsymbol{\theta}\mid\mathcal{D}_{\text{train}}^{\mathcal{T}_{t}})=\frac{1}{2P}\sum_{\mu=1}^{P}\big{[}f_{\boldsymbol{\theta}}(\boldsymbol{\xi}^{t,\mu})-\sigma^{t,\mu}\big{]}^{2},$ (2)
@@
The quality of $\boldsymbol{\theta}_{t}$ for solving task $\mathcal{T}_{t}$ can be evaluated by inspecting the validation loss $\ell(\boldsymbol{\theta}_{t}\mid\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}})$ using $\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\nu},\sigma^{t,\nu})\}_{\nu=1}^{V}$. For regression tasks, we can again use the MSE loss
$\ell(\boldsymbol{\theta}_{t}\mid\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}})=\frac{1}{2V}\sum_{\nu=1}^{P}\big{[}f_{\boldsymbol{\theta}_{t}}(\boldsymbol{\xi}^{t,\nu})-\sigma^{t,\nu}\big{]}^{2}.$ (3)
@@
Note that $\boldsymbol{\theta}_{t}$ is a function of $\boldsymbol{\theta}_{0}$ according to Eq. (1), which can be denoted as $\boldsymbol{\theta}_{t}(\boldsymbol{\theta}_{0})$. The meta-learner would like to update its parameter $\boldsymbol{\theta}_{0}$ so as to reduce the validation losses of different tasks
$\boldsymbol{\theta}_{0}^{\text{next}}$ $\leftarrow\boldsymbol{\theta}_{0}-\eta_{0}\nabla_{\boldsymbol{\theta}_{0}}\sum_{\mathcal{T}_{t}}\ell\big{(}\boldsymbol{\theta}_{t}(\boldsymbol{\theta}_{0})\mid\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}\big{)},$ (4)
@@
where $\eta_{0}$ is the learning rate for meta-training. After many episodes of meta-training using Eq. (4), we hopefully obtain a meta-learner which has pooled the knowledge of different tasks into the meta-learner’s parameters $\boldsymbol{\theta}_{0}$. The applicability of this meta-learner should be evaluated on an unseen task $\mathcal{T}_{*}$ which is assumed be drawn from $\mathcal{P}(\mathcal{T})$. Its performance can be assessed by the meta-generalization error $\epsilon^{\text{meta}}_{g}$, defined as
@@
$\epsilon^{\text{meta}}_{g}$ $=\big{\langle}\ell\big{(}\boldsymbol{\theta}_{*}(\boldsymbol{\theta}_{0})\mid\mathcal{D}_{\text{test}}^{\mathcal{T}_{*}}\big{)}\big{\rangle}_{\mathcal{T}_{*}\sim\mathcal{P}(\mathcal{T})},$ (5)
$\boldsymbol{\theta}_{*}(\boldsymbol{\theta}_{0})$ $:=\boldsymbol{\theta}_{0}-\eta_{*}\big{[}\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}\mid\mathcal{D}_{\text{train}}^{\mathcal{T}_{*}})\big{]}\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (6)
@@
where $\mathcal{D}_{\text{test}}^{\mathcal{T}_{*}}$ is a test dataset corresponding to the task $\mathcal{T}_{*}$. As such, the meta-generalization error $\epsilon^{\text{meta}}_{g}$ measures how good the meta-parameters $\boldsymbol{\theta}_{0}$ are in guiding the parameters $\boldsymbol{\theta}_{*}$ for solving a particular task $\mathcal{T}_{*}$ drawn from $\mathcal{P}(\mathcal{T})$.
@@
The MAML algorithm outlined above constitutes a bilevel optimization process, where Eq. (1) and Eq. (4) are referred to as the inner-loop update and the outer-loop update, respectively. In practice, the one-step inner-loop update in Eq. (1) has been generalized for better performance [29, 30], but the resulting algorithms are even more convoluted. The bilevel optimization structure in the original MAML setting has already involved second-order derivatives (e.g., in the outer-loop update), which presents challenges to computations in high dimensions. Approximations using only the first-order derivatives have been propose to reduce computations [31]. Another study observed that when using deep neural networks for solving computer vision tasks, the meta-learner tends to reuse the features for solving a new task, i.e., without much further tuning the representation [14]. Based on this empirical observation, the authors proposed the Almost-No-Inner-Loop (ANIL) algorithm, where only the parameters of the last layer of the neural networks are updated in the inner-loop optimization. In light of these earlier works, we focus on the First-Order ANIL (FO-ANIL) algorithm for simplification, where the algorithmic implementations will be described below.
@@ III Teacher-student Analysis
In this work, we extend the traditional teacher-student framework for supervised learning to the context of meta-learning. In the study of supervised learning, the core idea of the teacher-student analysis involves designing non-trivial teacher functions as the ground truths. The teacher functions generate input-output examples which are then presented to the student models for learning. It should be noted that the student models do not have access to the internal details of the teacher models, but only to the examples provided by them. In certain cases, the generalization ability of the student model can be expressed analytically through macroscopic order parameters that link the weights of the teacher and student models. This analytical approach enables us to conduct detailed theoretical examinations and gain deeper insights into the learning process [21, 20, 23].

### v2

@@ Abstract
Gradient-based meta-learning algorithms have gained popularity for their ability to train models on new tasks using limited data. Empirical observations indicate that such algorithms are able to learn a shared representation across tasks, which is regarded as a key factor in their success. However, the in-depth theoretical understanding of the learning dynamics and the origin of the shared representation remains underdeveloped. In this work, we investigate the meta-learning dynamics of nonlinear two-layer neural networks trained on streaming tasks in the teacher-student scenario. Through the lens of statistical physics analysis, we characterize the macroscopic behavior of the meta-training processes, the formation of the shared representation, and the generalization ability of the model on new tasks. The analysis also points to the importance of the choice of certain hyperparameters of the learning algorithms.
@@ I Introduction
Supervised learning has long been the dominant paradigm in machine learning and artificial intelligence, yielding substantial achievements across various fields [1]. The primary objective of supervised learning is to obtain a function mapping tailored to a specific task by training the machines on a set of input-output examples. However, one major limitation of this paradigm is that the trained machine is adapted to the specific task considered and is typically not suited for other tasks, even though they may share some common properties with the existing task. In extreme scenarios, it may require training a new model from scratch for each new task. This is not desirable if the supervised learning tasks require a large number of input-output examples, a common requirement for training large-scale deep neural networks [2, 3].
@@
To overcome this drawback, many alternative machine learning pipelines have been proposed, such as multitask learning, transfer learning, self-supervised learning, meta-learning, and so on [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A common strategy in the recent deep learning practice is to learn a useful representation of the data, which will become applicable to many different downstream tasks [6, 7, 8]. Alternatively, meta-learning typically aims at learning to how learn the algorithms for new tasks. Among various meta-learning approaches, the gradient-based meta-learning algorithms have received significant attention [11, 12]. A notable example is the model-agnostic meta-learning (MAML) algorithm [12]. It aims to learn a good initialization of the model parameters such that the model can be quickly adapted to new tasks starting from this set of initial model parameters; these initial model parameters are then meta-trained on many different tasks. As such, it involves two nested levels of optimization processes, both of which are implemented by gradient descent algorithms in MAML. Subsequent empirical work speculated that the success of MAML is also achieved by learning a shared representation across tasks and further proposed methods to simplify MAML [14].
@@
However, the nested optimization processes make it challenging to develop theoretical analyses for meta-learning algorithms like MAML. To facilitate the analysis, researchers often utilize simplified architectures, such as linear models [15, 16]. Notably, recent works investigated the learning dynamics of two-layer linear networks and proved that they can successfully learn the hidden representation under the learning dynamics with the population gradients in use [17, 18]. It is unknown whether similar results can be extended to more realistic settings of nonlinear neural networks and the learning dynamics with finite-sample gradients. In this paper, we aim to tackle the challenging scenarios with nonlinear network by employing tools from statistical physics, which has a longstanding tradition in the theoretical development of artificial neural networks [19, 20, 21], and has regained momentum in recent years [22, 23]. These methods have been recently applied to the multitask learning scenario [24, 25, 26, 27, 28] to study phenomena such as catastrophic forgetting and the effects of task similarity, as well as to propose novel learning algorithms; in these studies, the number of tasks is finite and there are abundant examples for each task. In this work, we focus on a scenario that is more typical in meta-learning, where the number of tasks is large, but the number of examples per task is much smaller than in single-task settings. We will further consider that the tasks arrive in a streaming, online fashion, which is similar to the usual continual learning setting. However, meta-learning approaches have greater emphasis on the ability to adapt to new tasks and require additional training steps for a new task, which differs from the usual continual learning setting.
@@ II The Problem and Algorithm of Meta-learning
The primary objective of meta-learning is to create a framework where a model can learn to adapt to some new task using a few examples. The model involves a meta-learner which is trained on different tasks in the meta-training phase, aiming to develop a meta-level knowledge that is useful for future tasks.
@@
In the meta-learning problem setup, there are typically many tasks available, which are denoted as $\{\mathcal{T}_{t}\}_{t=1,2,\dots}$ and assumed to be drawn from a certain distribution $\mathcal{P}(\mathcal{T})$. Each task $\mathcal{T}_{t}$ comes with its own training and validation datasets, denoted as $D_{\text{train}}^{\mathcal{T}_{t}}$ and $D_{\text{val}}^{\mathcal{T}_{t}}$, respectively. We focus on supervised learning tasks, so both the training and validation sets consist of input-output example pairs. Suppose a specific learner $\mathcal{L}_{t}$ is assigned to solve task $\mathcal{T}_{t}$. Its model weights are trained on the training set $D_{\text{train}}^{\mathcal{T}_{t}}$ and evaluated on the validation set $D_{\text{val}}^{\mathcal{T}_{t}}$. If the learner $\mathcal{L}_{t}$ inherits useful meta-level knowledge about the task $\mathcal{T}_{t}$ from an experienced meta-learner, it should require significantly fewer training examples than if it were learning from scratch. The challenge is how to effectively obtain such an experienced meta-learner. The model-agnostic meta-learning (MAML) algorithm addresses this by maintaining and adapting the meta-learner which serves as a valuable initialization (denoted as $\boldsymbol{\theta}_{0}$) for the model weights $\boldsymbol{\theta}_{t}$ of the learner $\mathcal{L}_{t}$. Since the learner $\mathcal{L}_{t}$ does not start from scratch, but has inherited the good initialization $\boldsymbol{\theta}_{0}$ from the meta-learner, it may require only a few steps of gradient-descent update by using the training set $D_{\text{train}}^{\mathcal{T}_{t}}$ in order to solve the task $\mathcal{T}_{t}$. The parameters $\boldsymbol{\theta}_{0}$ of the meta-learner are then updated also in a gradient-descent manner in order to reduce the validation error of the tasks that have shown up. Note that the validation error of task $\mathcal{T}_{t}$ is defined on the validation set $D_{\text{val}}^{\mathcal{T}_{t}}$.
@@
Formally, the learner $\mathcal{L}_{t}$ would like to learn a function $f_{\boldsymbol{\theta}_{t}}$ parameterized by $\boldsymbol{\theta}_{t}$ for solving task $\mathcal{T}_{t}$. It does so by starting from the meta-learner’s weights $\boldsymbol{\theta}_{0}$ and performing gradient descent using the training data $D_{\text{train}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\mu},\sigma^{t,\mu})\}_{\mu=1}^{P}$. In MAML, a one-step gradient descent is applied [12],
$\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{0}-\eta_{t}\big{[}\boldsymbol{\nabla}_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}\mid D_{\text{train}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (1)
@@
$\ell(\boldsymbol{\theta}\mid D_{\text{train}}^{\mathcal{T}_{t}})=\frac{1}{2P}\sum_{\mu=1}^{P}\big{[}f_{\boldsymbol{\theta}}(\boldsymbol{\xi}^{t,\mu})-\sigma^{t,\mu}\big{]}^{2},$ (2)
@@
The quality of $\boldsymbol{\theta}_{t}$ for solving task $\mathcal{T}_{t}$ can be evaluated by inspecting the validation loss $\ell(\boldsymbol{\theta}_{t}\mid D_{\text{val}}^{\mathcal{T}_{t}})$ using $D_{\text{val}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\nu},\sigma^{t,\nu})\}_{\nu=1}^{V}$. For regression tasks, we can again use the MSE loss
$\ell(\boldsymbol{\theta}_{t}\mid D_{\text{val}}^{\mathcal{T}_{t}})=\frac{1}{2V}\sum_{\nu=1}^{V}\big{[}f_{\boldsymbol{\theta}_{t}}(\boldsymbol{\xi}^{t,\nu})-\sigma^{t,\nu}\big{]}^{2}.$ (3)
@@
Note that $\boldsymbol{\theta}_{t}$ is a function of $\boldsymbol{\theta}_{0}$ according to Eq. (1), which can be denoted as $\boldsymbol{\theta}_{t}(\boldsymbol{\theta}_{0})$. The meta-learner would like to update its parameter $\boldsymbol{\theta}_{0}$ so as to reduce the validation losses of different tasks,
$\boldsymbol{\theta}_{0}^{\text{next}}$ $\leftarrow\boldsymbol{\theta}_{0}-\eta_{0}\nabla_{\boldsymbol{\theta}_{0}}\sum_{\mathcal{T}_{t}}\ell\big{(}\boldsymbol{\theta}_{t}(\boldsymbol{\theta}_{0})\mid D_{\text{val}}^{\mathcal{T}_{t}}\big{)},$ (4)
where $\eta_{0}$ is the learning rate for meta-training. Note that the update rule in Eq. (4) is designed to minimize the validation loss defined by the validation data sets $\{D_{\text{val}}^{\mathcal{T}_{t}}\}$. These validation datasets are included because the primary objective of the meta-learner is to guide task-specific learners to effectively generalize. Relying solely on the training data sets $\{D_{\text{train}}^{\mathcal{T}{t}}\}$ would not provide a reliable estimate of generalization ability.
@@
After many episodes of meta-training using Eq. (4), we hope to obtain a meta-learner which has pooled the knowledge of different tasks into the meta-learner’s parameters $\boldsymbol{\theta}_{0}$. The applicability of this meta-learner should be evaluated on an unseen task $\mathcal{T}_{*}$ which is assumed be drawn from $\mathcal{P}(\mathcal{T})$. Its performance can be assessed by the meta-generalization error $\epsilon^{\text{meta}}_{g}$, defined as
@@
$\epsilon^{\text{meta}}_{g}$ $=\big{\langle}\ell\big{(}\boldsymbol{\theta}_{*}(\boldsymbol{\theta}_{0})\mid D_{\text{test}}^{\mathcal{T}_{*}}\big{)}\big{\rangle}_{\mathcal{T}_{*}\sim\mathcal{P}(\mathcal{T})},$ (5)
$\boldsymbol{\theta}_{*}(\boldsymbol{\theta}_{0})$ $:=\boldsymbol{\theta}_{0}-\eta_{*}\big{[}\boldsymbol{\nabla}_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}\mid D_{\text{train}}^{\mathcal{T}_{*}})\big{]}\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (6)
@@
where $D_{\text{test}}^{\mathcal{T}_{*}}$ is a test dataset corresponding to the task $\mathcal{T}_{*}$. As such, the meta-generalization error $\epsilon^{\text{meta}}_{g}$ measures how good the meta-parameters $\boldsymbol{\theta}_{0}$ are in guiding the parameters $\boldsymbol{\theta}_{*}$ for solving a particular task $\mathcal{T}_{*}$ drawn from $\mathcal{P}(\mathcal{T})$.
@@
The MAML algorithm outlined above constitutes a bilevel optimization process, where Eq. (1) and Eq. (4) are referred to as the inner-loop update and the outer-loop update, respectively. In practice, the one-step inner-loop update in Eq. (1) has been generalized for better performance [29, 30], but the resulting algorithms are even more convoluted. While the MAML algorithm is conceptually simple, the bilevel optimization structure in the original MAML setting has already involved second-order derivatives (e.g., in the outer-loop update), which presents challenges to computations in high dimensions. Approximations using only the first-order derivatives have been propose to reduce computations [31]. Another study observed that when using deep neural networks for solving computer vision tasks, the meta-learner tends to reuse the features for solving a new task, i.e., without much further tuning the representation [14]. Based on this empirical observation, the authors proposed the almost-no-inner-loop (ANIL) algorithm, where only the parameters of the last layer of the neural networks are updated in the inner-loop optimization. In light of these earlier works, we focus on the first-order ANIL (FO-ANIL) algorithm for simplification, where the algorithmic implementations will be described below.
@@ III Teacher-student Analysis
In this work, we extend the traditional teacher-student framework for supervised learning to the context of meta-learning. In the study of supervised learning, the core idea of the teacher-student analysis involves designing nontrivial teacher functions as the ground truths. The teacher functions generate input-output examples which are then presented to the student models for learning. It should be noted that the student models do not have access to the internal details of the teacher models, but only to the examples provided by them. In certain cases, the generalization ability of the student model can be expressed analytically through macroscopic order parameters that link the weights of the teacher and student models. This analytical approach enables us to conduct detailed theoretical examinations and gain deeper insights into the learning process [21, 20, 23].